OpenStudy (anonymous):

Two tiny spheres of mass m = 9.80 mg carry charges of equal magnitude, 72.0 nC, but opposite sign. They are tied to the same ceiling hook by light strings of length 0.530 m. When a horizontal uniform electric field E that is directed to the left is turned on, the spheres hang at rest with the angle theta between the strings equal to 50.0

OpenStudy (anonymous):

hello are you trying to solve for the force F here?

OpenStudy (anonymous):

no i need the magnitude of the field

OpenStudy (anonymous):

\[F_q=kq^2/r^2\] \[r=2Lsin(\theta/2)\] force in the y \[Tcos(\theta/2)=mg\] force in the x \[F_E=F_q+mgtan(\theta/2)\] \[E=F_E/q\]

OpenStudy (anonymous):

This is what I have so far, and when I solve for it I'm getting an incorrect answer

OpenStudy (anonymous):

F=kQ1Q2/d^2, where k=constant=6.25x10^-9 Coulomb F=(6.25x10^-9)(72x10^-9)(72x10^9)/(.53)^2=1.15 x10^22 Newton

OpenStudy (anonymous):

for r I got .4479ish, .53 is L the length from the ball to the ceiling

OpenStudy (anonymous):